OpenStudy (anonymous):

Evaluate x3y * x2y2 Answer: a. x to the sixth, y squared b. 5x3y c. 6x2y d. x to the fifth, y cubed e. xy

OpenStudy (hba):

Why are you posting the question at more tabs ?

OpenStudy (hba):

Close this tab and come back :(

OpenStudy (anonymous):

well does x equal 2

OpenStudy (anonymous):

Ok... skittles hope that your your question is\[(x ^{3} \times y) \times (x ^{2} \times y ^{2})\]

OpenStudy (anonymous):

am I right?

OpenStudy (hba):

Yes it is exactly that lol @Rosh007

OpenStudy (hba):

By the way she cannot even see what it is @Rosh007

OpenStudy (anonymous):

why you guys have to be mean like that

OpenStudy (hba):

I am just telling him that you cannot see lol :D

OpenStudy (hba):